24-1: The net power produced by an air-standard Brayton cycle is 1500 kW. The pressure ratio for the cycle is 12 and the minimum and maximum temperatures are 20˚C and 1450˚C respectively. The isentropic efficiencies of the compressor and turbine are 0.75 and 0.80 respectively. Assume an ambient pressure of 100 kPa. Determine:

a. The thermal efficiency of the cycle

b. The mass flow rate of the air, in kg/s.

c. The heat added during the combustion process, in kW.

d. Allow the pressure ratio to vary and find the pressure ratio at which the cycle thermal efficiency is maximum.

e. 24-2: Repeat problem 24-1 using a cold air-standard basis with the specific heat values evaluated at 20˚C. In part d, compare the optimum pressure ratio to the theoretical optimal pressure ratio for the cold-air standard analysis defined by P_r,opt = (T3/T1)^[k/(2*(k-1))]
Answer #1

3 Gin data 4 (20 +273) kdvin T = 293K (iSo23 kin T123 K Ambient phebune (P)- 100 KA Ti T.= 293メ(12) C Scanned with CamScanne2 乡0.75-157591-293) CT 23) T. = 696.9 K -) = (12 .- T4 -8171s gsenわepic e易.con a もTJwa 723- 87.18 -.. T, = 0 22.35 K Let nご) +293) ISoo Sec 5.0216 x 1.ooS 123 696-9 5186.67 KT 2-3 Cyeta ett.cena, = Heat addod 500 5186.67 29 unal Cs Scanned with CaFa maximum wK atpt 293 22.2 = 58-74 Scanned with CamScanner